# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ V QGT determines asymptotic conversion rate
Importantly, if we regard $g\in G$ as an unknown parameter, this sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ achieve a conversion of a statistical model $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{g\in G}$ to another statistical model $\{\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\}_{g\in G}$ in the asymptotic limit. The asymptotic behavior of i.i.d. copies of parametric families of states has been studied in the context of QLAN [88, 89, 86, 87], which aims to approximate i.i.d. statistical models by a limit model, called the Gaussian shift model. In the following, after briefly reviewing QLAN for i.i.d. pure-state models in [86, 87], we point out that QGT is a crucial characteristic of their asymptotics. Based on this observation, we construct quantum channels asymptotically converting a pure state to another pure state in RTA, proving Proposition 6.
V.1 Quantum local asymptotic normality
Here, let us briefly review studies in QLAN [88, 89, 86, 87], which investigate the asymptotic properties of statistical models by associating them with their limit models.
We review particularly for QLAN on pure-state models [86, 87]. For a pure state $\psi$ and a set of Hermitian operators $\bm{X}=\{X_{i}\}_{i=1}^{m}$, we define a pure-state statistical model $\mathcal{U}_{\bm{\theta}}(\psi)$ with $\mathcal{U}_{\bm{\theta}}(\cdot)=U_{\bm{\theta}}(\cdot)U_{\bm{\theta}}^{\dagger}$ and $U_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}$, where $\bm{\theta}\cdot\bm{X}\coloneqq\sum_{i=1}^{m}\theta^{i}X_{i}$ for $\bm{\theta}\in\mathbb{R}^{m}$. To study its asymptotic property, we consider its i.i.d. copies of the state $(\mathcal{U}_{\bm{u}/\sqrt{N}}(\psi))^{\otimes N}$, where we introduced a local parameter $\bm{u}\coloneqq\sqrt{N}\bm{\theta}$. In the limit of $N\to\infty$, it is shown [86, 87] that this model is approximately interconvertible to the so-called Gaussian shift model with vanishing error.
[…]
$=0$ (44)
for $\epsilon\in(0,1/9)$.
V.2 QGT and asymptotic convertibility among i.i.d. statistical models
[…]
Importantly, the $(i,j)$ element of $C^{\dagger}C$ is given by
$\left(C^{\dagger}C\right)_{ij}$ $=\sum_{k=1}^{d-1}\braket{\psi|X_{i}|k}\braket{k|X_{j}|\psi}$
$=\braket{\psi|X_{i}(I-\psi)X_{j}|\psi},$ (45)
meaning that $C^{\dagger}C=\mathcal{Q}^{\psi}$, where $\mathcal{Q}^{\psi}$ is the QGT for $\mathcal{U}_{\bm{\theta}}(\psi)$. Therefore, if two pure-state statistical models have the same QGT, they are asymptotically interconvertible to each other since both of them can be reversibly converted to the same Gaussian shift model, shown in Eq. (44).
As a generalization of the above observation, we prove the following:
Lemma 8.
For sets of Hermitian operators $\bm{X}\coloneqq\{X_{i}\}_{i=1}^{m}$ and $\bm{X}^{\prime}\coloneqq\{X_{i}^{\prime}\}_{i=1}^{m}$, we define $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ and $\mathcal{U}^{\prime}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}$ for $\bm{\theta}\in\mathbb{R}^{m}$. For given two pure states $\psi$ and $\phi$, we define pure-state statistical models by $\mathcal{U}_{\bm{\theta}}(\psi)$ and $\mathcal{U}^{\prime}_{\bm{\theta}}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$, respectively. If $r>0$ satisfies $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$, then there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0$ (46)
for $\epsilon\in(0,1/9)$.
In the proof of Lemma 8, the argument is divided into two parts: one focuses on the case where the QGT is conserved, while the other addresses the case where the QGT is reduced. We first prove that the growth rate of i.i.d. copies can be set to one by adjusting the scaling of the parameters. Concretely, $\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}$ is shown to be asymptotically interconvertible to $\mathcal{U}^{\prime}_{\frac{\sqrt{r}\bm{u}}{\sqrt{N}}}(\phi)^{\otimes N}$. This asymptotic interconversion is achieved by combining the reversible conversion channels in QLAN, where the QGT is conserved. Then, we also prove that $\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)^{\otimes N}$ can be asymptotically converted into $\mathcal{U}^{\prime}_{\frac{\sqrt{r}\bm{u}}{\sqrt{N}}}(\phi)^{\otimes N}$ if the QGT with respect to $\bm{u}$ decreases in the sense of matrix inequality. The details of these lemmas and their proofs are provided in Appendix C.2.
V.3 Proof of achievability of the conversion rate
The process of transforming $\mathcal{U}_{g}(\psi)^{\otimes N}$ to $\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}$ with asymptotically vanishing error consists of two steps, which we may call the estimation and conversion steps. In the estimation step, we obtain a rough estimate $\hat{g}$ of the true value $g$ by consuming copies of $\mathcal{U}_{g}(\psi)$ sublinear in $N$. In the conversion step, based on the estimate $\hat{g}$, we convert the remaining copies of $\mathcal{U}_{g}(\psi)$ to i.i.d. copies of $\mathcal{U}_{g}^{\prime}(\phi)$ with quantum channels constructed by using Lemma 8.
For a detailed explanation, let us first introduce notations. Let $U$ and $U^{\prime}$ be (non-projective) unitary representations of a Lie group $G$ on the input and output systems. Elements in the neighborhood of the identity $e\in G$ can be parametrized as $g(\bm{\lambda})=e^{\mathrm{i}\sum_{\mu=1}^{\dim G}\lambda^{\mu}A_{\mu}}$ by using a basis $\{A_{\mu}\}_{\mu=1}^{\dim G}$ of the Lie algebra $\mathfrak{g}$. Define operators $X_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U(g(\bm{\lambda}))|_{\bm{\lambda}=\bm{0}}$ and $X^{\prime}_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U^{\prime}(g(\bm{\lambda}))|_{\bm{\lambda}=\bm{0}}$, which correspond to the images of $A_{\mu}$ by the Lie algebra representations induced from $U$ and $U^{\prime}$. At least locally, these operators satisfy $U(g(\bm{\lambda}))=e^{\mathrm{i}\sum_{\mu=1}^{\dim G}\lambda^{\mu}X_{\mu}}$ and $U^{\prime}(g(\bm{\lambda}))=e^{\mathrm{i}\sum_{\mu=1}^{\dim G}\lambda^{\mu}X_{\mu}^{\prime}}$. In this case, by slightly abusing our notation of $\mathcal{U}_{g}$ and $\mathcal{U}^{\prime}_{g}$, we also denote $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ and $\mathcal{U}_{\bm{\theta}}^{\prime}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}$ for $\bm{\theta}\in\mathbb{R}^{\dim G}$, implicitly indicating $\|\bm{\theta}\|$ is sufficiently small.
$G_{\mathrm{succ.}}^{(g,\delta)}\coloneqq\{\hat{g}\in G\mid\exists\bm{\theta},\,\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\bm{\theta}\|<\delta\}.$ (47)
We quantify the success probability in the worst case by
[…]
where $\mu_{G}$ is the Haar measure on the Lie group $G$, and $\chi_{A}$ denotes the indicator function of a set $A$. Note that the above success probability is well-defined since $G_{\mathrm{succ.}}^{(g,\delta)}$ is a measurable set, as shown in Appendix C.3.
Lemma 9.
$\lim_{N\to\infty}p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=1.$ (49)
We remark that $\|\bm{\theta}\|\sim N^{-1/2}$ for the error is the standard scaling in parameter estimation for i.i.d. models [90], including QLAN studies [88, 89, 86, 87]. The proof is detailed in Appendix C.4.
Lemma 10.
[…]
$\delta_{N}(g,\bm{u})\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}\left(\psi_{g}\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}\left(\phi_{g}\right)^{\otimes\lfloor rN\rfloor}\right)$ (50)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|\bm{u}\|<N^{\epsilon}}\delta_{N}(g,\bm{u})=0$ for $\epsilon\in(0,1/9)$.
This lemma is proven from Lemma 8 by using properties of a compact Lie group. The proof is detailed in Appendix D.
Proposition 6 is proven by combining Lemmas 9 and 10.
Proof of Proposition 6.
$\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N})$
$\coloneqq\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right).$ (51)
From the assumption $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, if parameters $\bm{\theta}$ satisfy $\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ for some $g,\hat{g}\in G$, then it also holds $\mathcal{U}^{\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}_{g}^{\prime}(\phi)$. Therefore, for $\hat{g}\in G_{\mathrm{succ.}}^{(g,\delta)}$ with $\delta>0$, we have
$\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\bm{\theta}}(\psi_{\hat{g}})^{\otimes N^{\prime}}\right),\mathcal{U}_{\bm{\theta}}^{\prime}(\phi_{\hat{g}})^{\otimes\lfloor rN^{\prime}\rfloor}\right),$ (52)
where we have defined $\psi_{\hat{g}}\coloneqq\mathcal{U}_{\hat{g}}(\psi)$ and $\phi_{\hat{g}}\coloneqq\mathcal{U}_{\hat{g}}^{\prime}(\phi)$, and $\mathrm{Fid}$ denotes the fidelity defined by $\mathrm{Fid}(\rho,\sigma)\coloneqq\left(\mathrm{Tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)\right)^{2}$ for states $\rho$ and $\sigma$. Since $N^{-1/2+\epsilon}<(N^{\prime})^{-1/2+\epsilon}$, we have
$f_{N}$
$\coloneqq\inf_{g\in G}\inf_{\hat{g}\in G_{\mathrm{succ.}}^{(g,N^{-1/2+\epsilon})}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\inf_{\hat{g}\in G}\inf_{\|\bm{u}\|<(N^{\prime})^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}(\psi_{\hat{g}})^{\otimes N^{\prime}}\right),\mathcal{U}_{\frac{\bm{u}}{\sqrt{N^{\prime}}}}^{\prime}(\phi_{\hat{g}})^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\left(1-\sup_{\hat{g}\in G}\sup_{\|\bm{u}\|<(N^{\prime})^{\epsilon}}\delta_{N^{\prime}}(\hat{g},\bm{u})\right)^{2}\xrightarrow{N\to\infty}1,$ (53)
where in the last line, we used the Fuchs–van de Graaf inequalities to relate the fidelity to the trance distance $\delta_{N}(g,\bm{u})$ defined in Eq. (50) in Lemma 10.
From Eqs. (49), (51), and (53), we get
$\inf_{g\in G}\mathrm{Fid}\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$=\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})$
[…]
$\geq p^{\mathrm{succ.}}(N^{-1/2+\epsilon})\times f_{N}$
$\xrightarrow{N\to\infty}1,$ (54)
implying that
From Lemma 7, this is equivalent to $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor(r-\delta)N\rfloor}\}_{N}$, which complete the proof of Proposition 6. ∎
The above proof of the direct part provides a clear understanding of the reason why the approximate asymptotic conversion rate diverges for finite groups. Since there is only a finite number of different elements in $\{\mathcal{U}_{g}(\psi)\mid g\in G\}$, it is possible to identify the state $\mathcal{U}_{g}(\psi)$ with an exponentially small failure probability using a state tomography protocol on $\mathcal{U}_{g}(\psi)^{\otimes N}$. Let $\hat{g}\in G$ denote the estimate of the true value $g$ obtained by this tomography protocol. Depending on the estimate $\hat{g}$, we can prepare an arbitrary number of copies of $\mathcal{U}^{\prime}_{\hat{g}}(\phi)$. The condition $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ implies that if $\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ is satisfied, then $\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}^{\prime}_{g}(\phi)$ also holds. Therefore, the approximate asymptotic conversion rate diverges since it is possible to create an arbitrary number of copies of $\mathcal{U}^{\prime}_{g}(\phi)$ from $\mathcal{U}_{g}(\psi)^{\otimes N}$ with asymptotically vanishing failure probability. We emphasize that the exact asymptotic conversion rate in [59] is not covered by this analysis because there always remains an error in the tomography process.
VI Circumventing constraint imposed by symmetry subgroup
When converting a pure state $\psi$ to another pure state $\phi$ in the asymptotic setup, the monotonicity of the QGTs yields an upper bound on possible conversion rate as shown in Sec. IV. However, the symmetry subgroups further restrict achievable conversion rate: the conversion rate always vanishes if $\mathrm{Sym}_{G}(\psi)\not\subset\mathrm{Sym}_{G}(\phi)$. In this section, we investigate ways to circumvent this restriction due to the symmetry subgroups by modifying the setup.
VI.1 Synchronization of asymmetry
$=\frac{V(\psi,H)}{V(\phi,H^{\prime})}+1,$ (58)
regardless of the periods of $\psi$ and $\phi$. Importantly, even when $R(\psi\to\phi)=0$, $\{\psi^{\otimes N}\otimes\phi^{\otimes N}\}_{N}$ can be approximately asymptotically convertible to $\{\phi^{\otimes\lfloor Nr^{*}\rfloor}\otimes\phi^{\otimes N}\}_{N}$ for any $r^{*}\in[0,V(\psi,H)/V(\phi,H^{\prime}))$. In this case, therefore, $\phi^{\otimes N}$ acts like a catalyst, enhancing the conversion rate from zero to $V(\psi,H)/V(\phi,H^{\prime})$ by bypassing the restriction on state convertibility imposed by their periods.
$R(\psi\otimes\phi\to\phi)$
VI.2 Sublinear additional resource
Theorem 11.
See Fig. 3 for a schematic picture of the setup of Theorem 11.
Proof of Theorem 11.
$\quad\geq\frac{M}{N}\bm{\gamma}^{\dagger}\mathcal{Q}^{\phi}\bm{\gamma}-\frac{M}{N}h(\epsilon)+\frac{1}{N}o\left(M\right),$ (62)
where $O^{\prime\prime}$ is defined by $O^{\prime\prime}\coloneqq\bm{\gamma}^{\dagger}\bm{X}^{\prime\prime}$ and $X^{\prime\prime}_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U^{\prime\prime}(g(\bm{\lambda}))|_{\bm{\lambda}=\bm{0}}$ by using the projective unitary representation of $G$ on the Hilbert space on which $\chi$ is defined. Since $\|\mathrm{i}[\chi,O^{\prime\prime}]\|_{f_{q},\chi}^{2}<\infty$ for any finite-dimensional system, the second term in the left-hand side vanishes in the limit of $N\to\infty$. By following the arguments from Eq. (38) to Eq. (40), we therefore get
Direct part: In the estimation step, we estimate $g\in G$ by consuming $(\mathcal{U}_{g}(\psi)\otimes\mathcal{U}_{g}^{\prime\prime}(\chi))^{\otimes\lceil N^{1-\epsilon}\rceil}$. The estimator $\hat{g}$ ensured to exist by Lemma 9 satisfies $((\mathcal{U}_{\bm{\theta}}\otimes\mathcal{U}^{\prime\prime}_{\bm{\theta}})\circ(\mathcal{U}_{\hat{g}}\otimes\mathcal{U}_{\hat{g}}^{\prime\prime}))(\psi\otimes\chi)=(\mathcal{U}_{g}\otimes\mathcal{U}^{\prime\prime}_{g})(\psi\otimes\chi)$, i.e.,
$\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)\,\land\,\mathcal{U}^{\prime\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime\prime}_{\hat{g}}(\chi)=\mathcal{U}^{\prime\prime}_{g}(\chi)$ (64)
with a failure probability that asymptotically vanishes as $N\to\infty$. From the assumption $\mathrm{Sym}_{G}(\chi)\subset\mathrm{Sym}_{G}(\phi)$, if $\bm{\theta}$ satisfies $\mathcal{U}^{\prime\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime\prime}_{\hat{g}}(\chi)=\mathcal{U}^{\prime\prime}_{g}(\chi)$, then $\mathcal{U}^{\prime}_{\bm{\theta}}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}^{\prime}_{g}(\phi)$ also holds, implying that Eq. (52) remains valid in this case.
We slightly modify the channel in Eq.(51) as
VII Mixed-state conversion
In the i.i.d. setting, the pure-state asymmetry conversion theory has been thoroughly developed for a compact Lie group in the preceding sections. This section extends the theory to mixed states, where the theoretical framework remains incomplete but is expected to serve as a foundation for further investigations.

### v4

@@ IV.2 Optimality
Importantly, if we regard $g\in G$ as an unknown parameter, this sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ achieves a conversion of a statistical model $\{\mathcal{U}_{g}(\psi)^{\otimes N}\}_{g\in G}$ to another statistical model $\{\mathcal{U}^{\prime}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\}_{g\in G}$ in the asymptotic limit.
We show that such an asymptotic transformation between statistical models is achieved by a two-step process: estimation and subsequent conversion. In the rest of this subsection, we assume that the unitary representations of $G$ are non-projective for simplicity.
Step 1 (Estimation): In this step, we obtain a rough estimation $\hat{g}$ of $g\in G$ using only a sublinear number of copies. More precisely, we apply an estimation procedure to $n=N^{1-\epsilon}$ copies of $\mathcal{U}_{g}(\psi)$ (for a fixed small $\epsilon$), thereby producing an estimate $\hat{g}$. The purpose of this step is not to learn $g$ precisely, but to localize it: with success probability tending to one as $N\to\infty$, the true state $\mathcal{U}_{g}(\psi)$ lies in a local neighborhood of $\mathcal{U}_{\hat{g}}(\psi)$, in the sense that, on the success event, there exists a small parameter $\theta\in\mathbb{R}^{\dim G}$ with $\|\theta\|=O(N^{-1/2+\epsilon})$ such that $\mathcal{U}_{g}(\psi)=e^{\mathrm{i}\theta^{\mu}X_{\mu}}\mathcal{U}_{\hat{g}}(\psi)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$.
After localization, we may treat the remaining copies (of order $N$) as a local i.i.d. model parameterized at the $N^{-1/2}$ scale. This is precisely the regime where quantum local asymptotic normality (QLAN) applies [70, 71, 72, 73], which approximates the properties of local i.i.d. models by Gaussian shift models. This will be exploited in the conversion step.
Notably, since this estimation step consumes only $n=N^{1-\epsilon}=o(N)$ copies, it incurs a sublinear overhead and hence does not change the achievable conversion rate in the $N\to\infty$ limit.
Step 2 (Conversion): In this step, conditioned on the estimate $\hat{g}$ obtained in Step 1, we convert the remaining $N^{\prime}=N-n$ copies of $\mathcal{U}_{g}(\psi)$ into copies of $\mathcal{U}^{\prime}_{g}(\phi)$. Since the success probability of Step 1 tends to one as $N\to\infty$, it suffices to describe the conversion on the success event, where the problem is reduced to a local i.i.d. model.
The key observation is that, in the local regime, QLAN translates pure-state unitary i.i.d. models by Gaussian shift models that are essentially characterized by the QGT. Importantly, from the additivity of the QGT, the Gaussian shift model associated with $\{\mathcal{U}_{g}(\phi)^{\otimes\lfloor rN\rfloor}\}_{g\in G}$ is governed by the scaled QGT, $r\,\mathcal{Q}^{\mathcal{U}_{g}(\phi)}$. Since the QGT quantifies the sensitivity of the state under infinitesimal transformations, the inequality
$\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\,\mathcal{Q}^{\mathcal{U}^{\prime}_{g}(\phi)}\qquad(\forall g\in G)$ (40)
can be interpreted as the condition that the Gaussian shift model associated with the input state is sufficiently “sensitive” to reproduce that for the output state at rate $r$ (uniformly in $g\in G$). This suggests that the conversion at a rate $r$ is possible if Eq. (40) is satisfied. This intuition is made rigorous in Appendix D by explicitly constructing conversion channels.
Here, the condition $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ is essential: estimation can only identify $g$ up to symmetries that leave $\psi$ invariant, and the inclusion guarantees that such residual ambiguity is also irrelevant for $\phi$. Hence, the conversion conditioned on $\hat{g}$ produces the correct target state $\mathcal{U}^{\prime}_{g}(\phi)$ whenever the localization in Step 1 succeeds.
Through the above two-step procedure, we construct a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ that achieves Eq. (39). Using Lemma 4, this establishes the opposite inequality of Eq. (38), i.e., the direct part of Eq. (9):
As detailed in Appendix D, our proof is constructive: we explicitly construct a sequence of $G$-covariant channels that achieves the asymptotic conversion. Also, we show that the conversion error decays polynomially in $N$.
V Mixed-state conversion
In the i.i.d. setting, the pure-state asymmetry conversion theory has been thoroughly developed for a compact Lie group in the preceding sections. This section partially extends the theory to mixed states. Although the theoretical framework remains incomplete, it is expected to serve as a foundation for further investigations.